OpenStudy (anonymous):

how do you find the value of "x" on a triangle on the Pythagorean Theorem? HELP!

OpenStudy (bahrom7893):

x^2+y^2=z^2 where x and y are the legs and z is the hypotenuse, so x = sqrt(z^2-y^2)

OpenStudy (anonymous):

okay can you help me with a problem and explain it please

OpenStudy (anonymous):

wwhere is the problem? in order to find x I need to see where they have placed it and what other information was provided

OpenStudy (anonymous):

pyth theorem is a^2 + b^2 = c^2 hypotheneuse is C

OpenStudy (anonymous):

so in your problem use a = x b = sqrt 10 c = sqrt 20 plug in what you have

OpenStudy (anonymous):